# Unifying Unsupervised Graph-Level Anomaly Detection and Out-of-Distribution Detection: A Benchmark

### v1

@@ 3.1 Benchmark Datasets
UB-GOLD includes 35 datasets in four types, each of which corresponds to either a GLAD or GLOD scenario categorized in Fig.1. Table 1 provides an overview of the benchmark datasets. For GLAD task, we consider two types of datasets with intrinsic anomaly and class-based anomaly to test the models’ performance of detecting anomalous graphs. For GLOD task, we consider datasets with inter-dataset shift and intra-dataset shift that assess the models’ ability to distinguish between ID and OOD samples. For more detailed information on the datasets, please see Appendix A.
Type I: Datasets with intrinsic anomaly. These datasets contain natural anomalies within chemical compounds, testing the robustness of GLAD methods. We use four datasets (i.e., Tox21_HSE, Tox21_MMP, Tox21_p53, and Tox21_PPAR-gamma) from the Tox21 challenge [29] involving molecules with unexpected biological activities.
Type II: Datasets with class-based anomaly. In these datasets, certain classes are designated as anomalies. We use 11 datasets (e.g., COLLAB, IMDB-BINARY, and ENZYMES) from the TU benchmark [30] where minority or distinct class samples are treated as anomalies.
Type III: Datasets with inter-dataset shift. We synthetic a benchmark dataset with inter-dataset shift by considering samples from one real-world dataset as ID and samples from another real-world dataset as OOD. For example, in “IMDB-MULTI→IMDB-BINARY”, graphs from IMDB-MULTI are ID, and graphs from IMDB-BINARY are OOD. The datasets belonging to the same domain and having close distribution shifts form a pair [9].
Type IV: Datasets with intra-dataset shift. Designed to assess GLOD methods under various types of intra-dataset shifts, these datasets are from graph OOD benchmarks, including GOOD [19] and DrugOOD [12]. Specifically, datasets with three kinds of domain shifts, i.e., assay shift, scaffold shift, and size shift, are considered.
@@ 3.2 Benchmark Algorithms
In UB-GOLD, we consider four groups of GLAD/GLOD methods for a comprehensive evaluation: 1) 2-Step methods, including graph kernel with detector (GK-D) and self-supervised learning (SSL) with detector (SSL-D), and 2) End-to-End methods, including GNN-based GLAD and GLOD methods. All methods are unsupervised, aligning with the primary focus of UB-GOLD to provide a comprehensive and fair evaluation framework. For a detailed summary of the models and further details, please refer to Table 2 and Appendix B.
@@
Graph kernel with detector. These methods follow a two-step process: obtaining graph embeddings using graph kernels and applying outlier detectors in the embedding space. For kernel methods, we consider Weisfeiler-Leman subtree kernel (WL) [31] and propagation kernel (PK) [21]. For detectors, we employ isolation forest (iF) [23] and one-class SVM (OCSVM, SVM for short) [22].
@@
GNN-based GLAD methods. We consider 6 SOTA methods for GLAD, including OCGIN [24], GLocalKD [3], OCGTL [7], SIGNET [8], GLADC [27], and CVTGAD [28]. These methods use different techniques, such as deep one-class classification, contrastive learning, and knowledge distillation, to detect anomalies in graph data.
Method Type | Category | Models and References |
2-Step | Graph kernel with detector [24] (GK-D) | PK-SVM [21; 22], PK-iF [23], WL-SVM [31] , WL-iF [31] |
Self-supervised learning with detector [9] (SSL-D) | IG-SVM[33], IG-iF[33], GCL-SVM[32], GCL-iF[32] | |
@@
End-to-End | Graph neural network-based GLAD | OCGIN[24], GLADC[27], GLocalKD [3], OCGTL [7], SIGNET [8], CVTGAD [28] |
@@
Graph neural network-based GLOD | GOOD-D [9], GraphDE [10] |
@@ 3.3 Evaluation Metrics and Implementation Details
Evaluation metrics. For comprehensive comparison, UB-GOLD utilizes three commonly used metrics for anomaly/OOD detection [34], i.e., AUROC, AUPRC, and FPR95. Higher AUROC and AUPRC values indicate better performance, while lower FPR95 values are preferable.
Data split. In our target scenarios (i.e., unsupervised GLAD/GLOD), all the samples in the training set are normal/ID, while the anomaly/OOD samples only occur in the testing set. In such an unsupervised case, the validation set with anomaly/OOD samples is usually unavailable during the training phase. Thus, following the implementation of OpenOOD [34], we divide the datasets into training and testing sets, without using a validation set. Specifically, we adopted the splits from [9] and [10], applying them to the benchmark datasets. Detailed splits are provided in Table 1.
@@
For more details related to the experimental setup in UB-GOLD, please refer to Appendix C.
@@ 4 Experimental Results & Discussion
GK-D (2-step) | SSL-D (2-step) | GNN-based GLAD (E2E) | GNN-based GLOD (E2E) | |||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | |
p53 | 49.17 | 54.05 | 57.69 | 54.40 | 68.11 | 60.85 | 68.61 | 64.60 | 68.35 | 65.43 | 67.58 | 68.10 | 65.82 | 69.40 | 67.82 | 62.59 |
HSE | 60.72 | 56.49 | 63.27 | 52.98 | 60.33 | 22.77 | 67.40 | 63.95 | 71.42 | 60.21 | 63.36 | 64.56 | 61.37 | 70.52 | 68.71 | 62.47 |
MMP | 51.03 | 49.95 | 55.50 | 51.98 | 57.72 | 52.58 | 69.91 | 71.31 | 69.37 | 68.12 | 67.51 | 71.23 | 70.03 | 70.58 | 71.41 | 60.12 |
PPAR | 53.74 | 48.42 | 57.76 | 49.45 | 61.78 | 63.22 | 68.37 | 69.88 | 67.75 | 65.29 | 66.43 | 68.88 | 69.43 | 68.83 | 68.21 | 66.31 |
COLLAB | 49.72 | 51.38 | 54.62 | 51.41 | 36.47 | 38.18 | 44.91 | 45.44 | 60.58 | 51.85 | 48.13 | 72.45 | 54.32 | 71.01 | 69.34 | 46.77 |
IMDB-B | 51.75 | 52.83 | 52.98 | 51.79 | 40.89 | 45.64 | 68.00 | 63.88 | 61.47 | 53.31 | 65.27 | 70.12 | 65.94 | 69.82 | 66.68 | 59.25 |
REDDIT-B | 48.36 | 46.19 | 49.50 | 49.84 | 60.32 | 52.51 | 84.49 | 82.64 | 82.10 | 80.32 | 89.92 | 85.24 | 78.87 | 87.43 | 89.43 | 63.42 |
ENZYMES | 52.45 | 49.82 | 53.75 | 51.03 | 60.97 | 53.94 | 62.73 | 63.09 | 62.44 | 61.75 | 63.59 | 63.12 | 63.44 | 68.56 | 64.58 | 52.10 |
PROTEINS | 49.43 | 61.24 | 53.85 | 65.75 | 61.15 | 52.78 | 72.61 | 72.60 | 76.46 | 77.29 | 72.89 | 75.86 | 77.43 | 76.49 | 76.02 | 68.81 |
DD | 47.69 | 75.29 | 47.98 | 70.49 | 70.33 | 42.67 | 76.43 | 65.41 | 79.08 | 80.76 | 77.76 | 74.53 | 76.54 | 78.84 | 79.91 | 60.49 |
BZR | 46.67 | 59.08 | 51.16 | 51.71 | 41.50 | 45.42 | 68.93 | 67.81 | 69.13 | 68.55 | 51.89 | 80.79 | 68.23 | 77.69 | 73.28 | 65.94 |
AIDS | 50.93 | 52.01 | 52.56 | 61.42 | 87.20 | 97.96 | 95.44 | 98.80 | 96.89 | 96.93 | 99.36 | 97.60 | 98.02 | 99.21 | 97.10 | 70.82 |
COX2 | 52.15 | 52.48 | 53.34 | 49.56 | 49.11 | 48.61 | 59.68 | 59.38 | 57.81 | 58.93 | 59.81 | 72.35 | 64.13 | 64.36 | 63.19 | 54.73 |
NCI1 | 51.39 | 50.22 | 54.18 | 50.41 | 45.11 | 61.88 | 43.33 | 46.44 | 69.46 | 65.29 | 75.75 | 74.32 | 68.32 | 69.13 | 61.58 | 58.74 |
DHFR | 48.31 | 52.79 | 50.30 | 51.64 | 45.58 | 63.15 | 58.21 | 57.01 | 61.09 | 61.79 | 59.82 | 72.87 | 61.25 | 63.23 | 64.48 | 53.23 |
IM→IB | 49.80 | 51.23 | 53.45 | 53.03 | 56.26 | 51.32 | 74.45 | 78.62 | 80.98 | 81.25 | 66.73 | 71.10 | 78.28 | 80.23 | 80.94 | 52.67 |
EN→PR | 52.53 | 53.36 | 53.92 | 51.90 | 46.01 | 33.52 | 59.76 | 63.23 | 61.77 | 59.36 | 67.18 | 62.42 | 56.95 | 64.31 | 63.84 | 54.48 |
AI→DH | 51.18 | 51.69 | 52.28 | 50.95 | 84.33 | 83.27 | 97.11 | 98.48 | 95.68 | 94.33 | 98.95 | 96.82 | 95.42 | 99.10 | 99.27 | 94.58 |
BZ→CO | 43.34 | 52.43 | 49.76 | 52.16 | 64.29 | 64.65 | 78.98 | 76.01 | 87.27 | 80.55 | 81.86 | 89.11 | 83.21 | 96.32 | 95.16 | 65.26 |
ES→MU | 52.99 | 52.63 | 52.13 | 52.28 | 58.12 | 51.57 | 78.66 | 79.66 | 86.70 | 90.55 | 88.32 | 91.43 | 89.30 | 92.41 | 91.98 | 75.65 |
TO→SI | 53.73 | 51.87 | 53.50 | 52.25 | 64.32 | 66.53 | 66.85 | 64.85 | 67.29 | 69.80 | 68.91 | 66.72 | 72.51 | 68.24 | 66.70 | 72.34 |
BB→BA | 54.15 | 53.11 | 54.62 | 53.48 | 63.27 | 32.37 | 69.18 | 67.33 | 78.83 | 77.69 | 78.93 | 89.88 | 79.07 | 80.17 | 81.44 | 55.69 |
PT→MU | 51.52 | 55.87 | 54.03 | 52.12 | 55.88 | 53.78 | 78.10 | 79.74 | 79.27 | 77.54 | 62.51 | 84.63 | 80.12 | 79.44 | 82.05 | 58.28 |
FS→TC | 50.06 | 54.76 | 51.98 | 53.24 | 44.98 | 49.57 | 67.05 | 66.01 | 66.98 | 68.92 | 64.38 | 78.12 | 67.32 | 69.89 | 71.58 | 60.12 |
CL→LI | 50.85 | 51.74 | 52.66 | 51.54 | 55.62 | 56.45 | 59.65 | 54.17 | 61.21 | 58.31 | 59.30 | 72.15 | 63.42 | 70.21 | 69.28 | 50.79 |
HIV-Size | 48.94 | 49.96 | 66.11 | 45.10 | 31.39 | 32.67 | 26.73 | 35.80 | 38.55 | 42.94 | 96.34 | 91.86 | 47.56 | 56.23 | 74.12 | 68.31 |
ZINC-Size | 48.66 | 50.12 | 50.58 | 48.96 | 53.07 | 53.47 | 52.61 | 53.46 | 54.22 | 54.21 | 59.41 | 57.29 | 52.53 | 58.78 | 68.43 | 55.63 |
HIV-Scaffold | 49.36 | 48.43 | 44.72 | 54.57 | 58.78 | 58.74 | 61.00 | 59.26 | 56.82 | 54.38 | 58.05 | 70.93 | 63.23 | 59.49 | 62.28 | 53.48 |
ZINC-Scaffold | 51.12 | 46.66 | 51.17 | 53.28 | 54.77 | 55.17 | 54.04 | 55.80 | 51.87 | 50.12 | 56.79 | 59.24 | 55.79 | 55.73 | 56.39 | 55.62 |
IC50-Size | 67.08 | 59.35 | 90.76 | 52.49 | 32.61 | 36.15 | 24.44 | 30.96 | 42.58 | 41.29 | 97.36 | 81.43 | 39.63 | 50.37 | 50.71 | 45.24 |
EC50-Size | 70.50 | 59.33 | 92.29 | 49.36 | 29.71 | 32.60 | 22.83 | 27.68 | 41.37 | 39.42 | 97.74 | 77.12 | 39.23 | 55.84 | 56.58 | 59.49 |
IC50-Scaffold | 66.33 | 60.95 | 88.49 | 58.56 | 36.96 | 38.67 | 34.60 | 38.82 | 44.56 | 42.97 | 96.00 | 77.58 | 38.43 | 57.96 | 56.62 | 59.68 |
EC50-Scaffold | 64.95 | 62.78 | 86.03 | 55.27 | 27.88 | 30.36 | 31.28 | 32.35 | 51.92 | 48.43 | 94.18 | 74.65 | 40.98 | 66.52 | 58.29 | 54.24 |
IC50-Assay | 54.47 | 53.13 | 59.18 | 52.11 | 51.23 | 51.42 | 51.15 | 51.93 | 52.61 | 49.87 | 68.78 | 65.99 | 52.12 | 54.25 | 53.74 | 57.11 |
EC50-Assay | 49.08 | 46.66 | 48.43 | 45.32 | 47.80 | 47.81 | 47.80 | 46.02 | 56.11 | 52.44 | 69.31 | 82.97 | 54.34 | 66.71 | 65.57 | 58.96 |
Avg. AUROC | 52.69 | 53.67 | 57.56 | 52.91 | 52.11 | 50.35 | 61.29 | 61.50 | 66.00 | 64.29 | 73.14 | 75.86 | 65.50 | 71.07 | 71.05 | 60.38 |
Avg. Rank | 12.09 | 11.94 | 10.46 | 11.77 | 12.06 | 11.65 | 8.14 | 8.31 | 6.69 | 8.23 | 4.91 | 3.34 | 7.14 | 4.51 | 4.54 | 10.14 |
@@
In this section, we introduce the experimental setup and discuss the experimental results in UB-GOLD benchmark. Specifically, we aim to answer the following research questions: ⚫ RQ1 (Effectiveness): How do different GLAD and GLOD methods perform under various anomaly scenarios and distribution shifts? (Sec. 4.1) ⚫ RQ2 (Generalizability): How effective are GLAD and GLOD methods in detecting near-OOD and far-OOD graph samples? (Sec. 4.2) ⚫ RQ3 (Robustness): How does the inclusion of OOD samples in the ID training set affect the robustness of GLAD and GLOD methods? (Sec. 4.3) ⚫ RQ4 (Efficiency): Are the GLAD and GLOD methods efficient in terms of run time and memory usage? (Sec. 4.4)

### v2

@@ 3.1 Benchmark Datasets
UB-GOLD includes 35 datasets in four types, each of which corresponds to either a GLAD or GLOD scenario categorized in Fig.1. Table 1 provides an overview of the benchmark datasets. For GLAD task, we consider two types of datasets with intrinsic anomaly and class-based anomaly to test the models’ performance of detecting anomalous graphs. For GLOD task, we consider datasets with inter-dataset shift and intra-dataset shift that assess the models’ ability to distinguish between ID and OOD samples. For more detailed information on the datasets, please see Appendix B.
Type I: Datasets with intrinsic anomaly. These datasets contain natural anomalies within chemical compounds, testing the robustness of GLAD methods. We use four datasets (i.e., Tox21_HSE, Tox21_MMP, Tox21_p53, and Tox21_PPAR-gamma) from the Tox21 challenge (Abdelaziz et al., 2016) involving molecules with unexpected biological activities.
Type II: Datasets with class-based anomaly. In these datasets, certain classes are designated as anomalies. We use 11 datasets (e.g., COLLAB, IMDB-BINARY, and ENZYMES) from the TU benchmark (Morris et al., 2020) where minority or distinct class samples are treated as anomalies.
Type III: Datasets with inter-dataset shift. We synthetic a benchmark dataset with inter-dataset shift by considering samples from one real-world dataset as ID and samples from another real-world dataset as OOD. For example, in “IMDB-MULTI→IMDB-BINARY”, graphs from IMDB-MULTI are ID, and graphs from IMDB-BINARY are OOD. The datasets belonging to the same domain and having close distribution shifts form a pair (Liu et al., 2023a).
Type IV: Datasets with intra-dataset shift. Designed to assess GLOD methods under various types of intra-dataset shifts, these datasets are from graph OOD benchmarks, including GOOD (Gui et al., 2022) and DrugOOD (Ji et al., 2023). Specifically, datasets with three kinds of domain shifts, i.e., assay shift, scaffold shift, and size shift, are considered.
@@ 3.2 Benchmark Algorithms
In UB-GOLD, we consider four groups of GLAD/GLOD methods for a comprehensive evaluation: 1) two-step methods, including graph kernel with detector (GK-D) and self-supervised learning (SSL) with detector (SSL-D), and 2) End-to-End methods, including GNN-based GLAD and GLOD methods. For the end-to-end methods, we further divided them into 4 categories from the technical perspective, including one-class classification-based, graph reconstruction-based, contrastive learning-based, and well-trained GNN-based. All methods are unsupervised, aligning with the primary focus of UB-GOLD to provide a comprehensive and fair evaluation framework. For a detailed summary of the models and further details, please refer to Table 2 and Appendix C.
@@
Graph kernel with detector. These methods follow a two-step process: obtaining graph embeddings using graph kernels and applying outlier detectors in the embedding space. For kernel methods, we consider Weisfeiler-Leman subtree kernel (WL) (Li et al., 2016) and propagation kernel (PK) (Neumann et al., 2016). For detectors, we employ isolation forest (iF) (Liu et al., 2008) and one-class SVM (OCSVM, SVM for short) (Amer et al., 2013).
SSL with detector. These approaches also follow a two-step process but use SSL methods to obtain graph embeddings. We consider two SSL methods, GraphCL (GCL for short) (You et al., 2020) and InfoGraph (IG for short) (Sun et al., 2020) to generate embeddings and use iF (Liu et al., 2008) and SVM (Amer et al., 2013) as detectors.
@@
GNN-based GLAD methods. We consider 6 SOTA methods for GLAD, including OCGIN (Zhao & Akoglu, 2023), GLocalKD (Ma et al., 2022), OCGTL (Qiu et al., 2022), SIGNET (Liu et al., 2023b), GLADC (Luo et al., 2022), and CVTGAD (Li et al., 2023). These methods use different techniques, such as deep one-class classification, contrastive learning, and knowledge distillation, to detect anomalies in graph data.
GNN-based GLOD methods. We involve four representatives of unsupervised GLOD methods, i.e., GOOD-D (Liu et al., 2023a), GraphDE (Li et al., 2022), AAGOD Guo et al., 2023, and GOODAT (Wang et al., 2024a) for comparison. GOOD-D is a contrastive learning-based approach, while GraphDE is a generative model-based approach. Both AAGOD and GOODAT operate on well-trained GNNs111In this context, “well-trained GNNs” refers to graph neural networks that have been pre-trained using a self-supervised or task-specific pretraining strategy, incorporating domain knowledge to effectively learn graph representations. These models are then fine-tuned for the specific GLAD or GLOD tasks.; AAGOD follows a data-centric approach, and GOODAT focuses on test-time OOD detection.
Method Type | Category | Models |
Two-Step | Graph kernel with detector (GK-D) | PK-SVM, PK-iF, WL-SVM, WL-iF |
Self-supervised learning with detector (SSL-D) | IG-SVM, IG-iF, GCL-SVM, GCL-iF | |
@@
End-to-End | Graph neural network-based GLAD | OCGIN, GLADC, GLocalKD, OCGTL, SIGNET, CVTGAD |
@@
Graph neural network-based GLOD | GOOD-D, GraphDE, AAGOD, GOODAT |
UB-GOLD: A Unified Benchmark Library for Unsupervised GLAD and GLOD. We offer our developed benchmark library UB-GOLD, as illustrated in Fig. 2. Specifically, the benchmark starts with a Benchmark Datasets module, which includes four types of datasets: intrinsic anomaly, class-based anomaly, inter-dataset shift, and intra-dataset shift. Each dataset type is designed to cover both GLAD and GLOD scenarios, ensuring broad applicability across different graph-level anomaly and OOD detection tasks. The Benchmark Tasks module prepares the settings (e.g., datasets splits and pre-processing) for different benchmarking tasks of anomaly and OOD detection. It supports three main tasks: general anomaly/OOD detection, near-far OOD detection, and perturbation-based detection, enabling models to handle different levels of OOD difficulty and assess robustness to data perturbations. This step ensures the data is well-prepared for downstream tasks. Next, in the Method module, UB-GOLD includes both two-step methods and end-to-end methods. Specifically, we categorize end-to-end GLAD and GLOD methods into four technical groups: one-class classification, graph reconstruction, well-trained GNNs, and contrastive learning. Finally, the Evaluation module ensures a comprehensive and fair assessment, offering Comprehensive Metrics, Robustness under Perturbation, Generalization on Near-OOD and Far-OOD, and OOD Judge Score Distribution Analysis to measure how well models distinguish between ID and OOD samples.
@@ 3.3 Evaluation Metrics and Implementation Details
Evaluation metrics. For comprehensive comparison, UB-GOLD utilizes three commonly used metrics for anomaly/OOD detection (Zhang et al., 2023), i.e., AUROC, AUPRC, and FPR95. Higher AUROC and AUPRC values indicate better performance, while lower FPR95 values are preferable.
Data split. In our target scenarios (i.e., unsupervised GLAD/GLOD), all the samples in the training set are normal/ID, while the anomaly/OOD samples only occur in the testing set. In such an unsupervised case, the validation set with anomaly/OOD samples is usually unavailable during the training phase. Thus, following the implementation of OpenOOD (Zhang et al., 2023), we divide the datasets into training and testing sets, without using a validation set. Specifically, we adopted the splits from (Liu et al., 2023a) and (Li et al., 2022), applying them to the benchmark datasets. Detailed splits are provided in Table 1.
@@
For more details related to the experimental setup in UB-GOLD, please refer to Appendix D.
@@ 4 Experimental Results & Discussion
GK-D (two-step) | SSL-D (two-step) | GNN-based GLAD (end-to-end) | GNN-based GLOD (end-to-end) | |||||||||||||||
@@
PK-SVM | PK-iF | WL-SVM | WL-iF | IG-SVM | IG-iF | GCL-SVM | GCL-iF | OCGIN | GLocalKD | OCGTL | SIGNET | GLADC | CVTGAD | GOOD-D | GraphDE | AAGOD | GOODAT | |
p53 | 49.17 | 54.05 | 57.69 | 54.40 | 68.11 | 60.85 | 68.61 | 64.60 | 68.35 | 65.43 | 67.58 | 68.10 | 65.82 | 69.40 | 67.82 | 62.59 | 50.12 | 61.71 |
HSE | 60.72 | 56.49 | 63.27 | 52.98 | 60.33 | 22.77 | 67.40 | 63.95 | 71.42 | 60.21 | 63.36 | 64.56 | 61.37 | 70.52 | 68.71 | 62.47 | 54.60 | 61.99 |
MMP | 51.03 | 49.95 | 55.50 | 51.98 | 57.72 | 52.58 | 69.91 | 71.31 | 69.37 | 68.12 | 67.51 | 71.23 | 70.03 | 70.58 | 71.41 | 60.12 | 62.92 | 67.49 |
PPAR | 53.74 | 48.42 | 57.76 | 49.45 | 61.78 | 63.22 | 68.37 | 69.88 | 67.75 | 65.29 | 66.43 | 68.88 | 69.43 | 68.83 | 68.21 | 66.31 | 57.90 | 66.95 |
COLLAB | 49.72 | 51.38 | 54.62 | 51.41 | 36.47 | 38.18 | 44.91 | 45.44 | 60.58 | 51.85 | 48.13 | 72.45 | 54.32 | 71.01 | 69.34 | 46.77 | 50.14 | 44.91 |
IMDB-B | 51.75 | 52.83 | 52.98 | 51.79 | 40.89 | 45.64 | 68.00 | 63.88 | 61.47 | 53.31 | 65.27 | 70.12 | 65.94 | 69.82 | 66.68 | 59.25 | 58.43 | 65.46 |
REDDIT-B | 48.36 | 46.19 | 49.50 | 49.84 | 60.32 | 52.51 | 84.49 | 82.64 | 82.10 | 80.32 | 89.92 | 85.24 | 78.87 | 87.43 | 89.43 | 63.42 | 68.78 | 80.31 |
ENZYMES | 52.45 | 49.82 | 53.75 | 51.03 | 60.97 | 53.94 | 62.73 | 63.09 | 62.44 | 61.75 | 63.59 | 63.12 | 63.44 | 68.56 | 64.58 | 52.10 | 58.70 | 52.33 |
PROTEINS | 49.43 | 61.24 | 53.85 | 65.75 | 61.15 | 52.78 | 72.61 | 72.60 | 76.46 | 77.29 | 72.89 | 75.86 | 77.43 | 76.49 | 76.02 | 68.81 | 75.04 | 75.92 |
DD | 47.69 | 75.29 | 47.98 | 70.49 | 70.33 | 42.67 | 76.43 | 65.41 | 79.08 | 80.76 | 77.76 | 74.53 | 76.54 | 78.84 | 79.91 | 60.49 | 74.00 | 77.62 |
BZR | 46.67 | 59.08 | 51.16 | 51.71 | 41.50 | 45.42 | 68.93 | 67.81 | 69.13 | 68.55 | 51.89 | 80.79 | 68.23 | 77.69 | 73.28 | 65.94 | 64.52 | 64.77 |
AIDS | 50.93 | 52.01 | 52.56 | 61.42 | 87.20 | 97.96 | 95.44 | 98.80 | 96.89 | 96.93 | 99.36 | 97.60 | 98.02 | 99.21 | 97.10 | 70.82 | 86.64 | 98.82 |
COX2 | 52.15 | 52.48 | 53.34 | 49.56 | 49.11 | 48.61 | 59.68 | 59.38 | 57.81 | 58.93 | 59.81 | 72.35 | 64.13 | 64.36 | 63.19 | 54.73 | 51.86 | 59.99 |
NCI1 | 51.39 | 50.22 | 54.18 | 50.41 | 45.11 | 61.88 | 43.33 | 46.44 | 69.46 | 65.29 | 75.75 | 74.32 | 68.32 | 69.13 | 61.58 | 58.74 | 49.94 | 45.96 |
DHFR | 48.31 | 52.79 | 50.30 | 51.64 | 45.58 | 63.15 | 58.21 | 57.01 | 61.09 | 61.79 | 59.82 | 72.87 | 61.25 | 63.23 | 64.48 | 53.23 | 63.93 | 61.52 |
IM→IB | 49.80 | 51.23 | 53.45 | 53.03 | 56.26 | 51.32 | 74.45 | 78.62 | 80.98 | 81.25 | 66.73 | 71.10 | 78.28 | 80.23 | 80.94 | 52.67 | 82.17 | 77.66 |
EN→PR | 52.53 | 53.36 | 53.92 | 51.90 | 46.01 | 33.52 | 59.76 | 63.23 | 61.77 | 59.36 | 67.18 | 62.42 | 56.95 | 64.31 | 63.84 | 54.48 | 50.17 | 64.61 |
AI→DH | 51.18 | 51.69 | 52.28 | 50.95 | 84.33 | 83.27 | 97.11 | 98.48 | 95.68 | 94.33 | 98.95 | 96.82 | 95.42 | 99.10 | 99.27 | 94.58 | 94.90 | 93.95 |
BZ→CO | 43.34 | 52.43 | 49.76 | 52.16 | 64.29 | 64.65 | 78.98 | 76.01 | 87.27 | 80.55 | 81.86 | 89.11 | 83.21 | 96.32 | 95.16 | 65.26 | 77.44 | 80.97 |
ES→MU | 52.99 | 52.63 | 52.13 | 52.28 | 58.12 | 51.57 | 78.66 | 79.66 | 86.70 | 90.55 | 88.32 | 91.43 | 89.30 | 92.41 | 91.98 | 75.65 | 91.91 | 83.92 |
TO→SI | 53.73 | 51.87 | 53.50 | 52.25 | 64.32 | 66.53 | 66.85 | 64.85 | 67.29 | 69.80 | 68.91 | 66.72 | 72.51 | 68.24 | 66.70 | 72.34 | 66.90 | 67.21 |
BB→BA | 54.15 | 53.11 | 54.62 | 53.48 | 63.27 | 32.37 | 69.18 | 67.33 | 78.83 | 77.69 | 78.93 | 89.88 | 79.07 | 80.17 | 81.44 | 55.69 | 72.00 | 75.94 |
PT→MU | 51.52 | 55.87 | 54.03 | 52.12 | 55.88 | 53.78 | 78.10 | 79.74 | 79.27 | 77.54 | 62.51 | 84.63 | 80.12 | 79.44 | 82.05 | 58.28 | 67.65 | 80.42 |
FS→TC | 50.06 | 54.76 | 51.98 | 53.24 | 44.98 | 49.57 | 67.05 | 66.01 | 66.98 | 68.92 | 64.38 | 78.12 | 67.32 | 69.89 | 71.58 | 60.12 | 67.65 | 67.09 |
CL→LI | 50.85 | 51.74 | 52.66 | 51.54 | 55.62 | 56.45 | 59.65 | 54.17 | 61.21 | 58.31 | 59.30 | 72.15 | 63.42 | 70.21 | 69.28 | 50.79 | 53.08 | 60.93 |
HIV-Size | 48.94 | 49.96 | 66.11 | 45.10 | 31.39 | 32.67 | 26.73 | 35.80 | 38.55 | 42.94 | 96.34 | 91.86 | 47.56 | 56.23 | 74.12 | 68.31 | 41.83 | 29.21 |
ZINC-Size | 48.66 | 50.12 | 50.58 | 48.96 | 53.07 | 53.47 | 52.61 | 53.46 | 54.22 | 54.21 | 59.41 | 57.29 | 52.53 | 58.78 | 68.43 | 55.63 | 56.56 | 52.51 |
HIV-Scaffold | 49.36 | 48.43 | 44.72 | 54.57 | 58.78 | 58.74 | 61.00 | 59.26 | 56.82 | 54.38 | 58.05 | 70.93 | 63.23 | 59.49 | 62.28 | 53.48 | 60.75 | 55.75 |
ZINC-Scaffold | 51.12 | 46.66 | 51.17 | 53.28 | 54.77 | 55.17 | 54.04 | 55.80 | 51.87 | 50.12 | 56.79 | 59.24 | 55.79 | 55.73 | 56.39 | 55.62 | 52.06 | 48.68 |
IC50-Size | 67.08 | 59.35 | 90.76 | 52.49 | 32.61 | 36.15 | 24.44 | 30.96 | 42.58 | 41.29 | 97.36 | 81.43 | 39.63 | 50.37 | 50.71 | 45.24 | 40.57 | 27.78 |
EC50-Size | 70.50 | 59.33 | 92.29 | 49.36 | 29.71 | 32.60 | 22.83 | 27.68 | 41.37 | 39.42 | 97.74 | 77.12 | 39.23 | 55.84 | 56.58 | 59.49 | 45.92 | 32.97 |
IC50-Scaffold | 66.33 | 60.95 | 88.49 | 58.56 | 36.96 | 38.67 | 34.60 | 38.82 | 44.56 | 42.97 | 96.00 | 77.58 | 38.43 | 57.96 | 56.62 | 59.68 | 49.79 | 36.34 |
EC50-Scaffold | 64.95 | 62.78 | 86.03 | 55.27 | 27.88 | 30.36 | 31.28 | 32.35 | 51.92 | 48.43 | 94.18 | 74.65 | 40.98 | 66.52 | 58.29 | 54.24 | 45.85 | 41.86 |
IC50-Assay | 54.47 | 53.13 | 59.18 | 52.11 | 51.23 | 51.42 | 51.15 | 51.93 | 52.61 | 49.87 | 68.78 | 65.99 | 52.12 | 54.25 | 53.74 | 57.11 | 48.74 | 45.34 |
EC50-Assay | 49.08 | 46.66 | 48.43 | 45.32 | 47.80 | 47.81 | 47.80 | 46.02 | 56.11 | 52.44 | 69.31 | 82.97 | 54.34 | 66.71 | 65.57 | 58.96 | 60.42 | 55.79 |
Avg. AUROC | 52.69 | 53.67 | 57.56 | 52.91 | 52.11 | 50.35 | 61.29 | 61.50 | 66.00 | 64.29 | 73.14 | 75.86 | 65.50 | 71.07 | 71.05 | 60.38 | 61.54 | 61.91 |
Avg. Rank | 13.80 | 13.54 | 12.03 | 13.86 | 14.00 | 13.83 | 10.83 | 10.63 | 7.26 | 8.71 | 5.43 | 3.37 | 7.03 | 3.69 | 4.14 | 10.58 | 9.89 | 9.43 |
@@
In this section, we introduce the experimental setup and discuss the experimental results in UB-GOLD benchmark. Specifically, we aim to answer the following research questions: ⚫ RQ1 (Effectiveness): How do different GLAD and GLOD methods perform under various anomaly scenarios and distribution shifts? (Sec. 4.1) ⚫ RQ2 (OOD sensitivity spectrum): How effective are GLAD and GLOD methods in detecting near-OOD and far-OOD graph samples? (Sec. 4.2) ⚫ RQ3 (Robustness): How does the inclusion of OOD samples in the ID training set affect the robustness of GLAD and GLOD methods? (Sec. 4.3) ⚫ RQ4 (Efficiency): Are the GLAD and GLOD methods efficient in terms of run time and memory usage? (Sec. 4.4)
These research questions are designed to comprehensively evaluate the performance, OOD sensitivity spectrum, robustness, and efficiency of GLAD and GLOD methods. By examining these aspects, we aim to provide a thorough understanding of their strengths and limitations across different scenarios. For more detailed experimental descriptions and settings, please refer to Appendix F.